# Software Performance Engineering for Foundation Model-Powered Software (FMware)

### v1

@@ III-A Challenge 1: The complexity of creating high-performance cognitive architectures
The first step in developing FMware is to create an appropriate cognitive architecture. A cognitive architecture defines how different AI components interact and reason together to achieve desired outcomes. This architecture complements the classical software architecture, detailing how AI reasoning and results are delivered through traditional software components like regular and/or vector databases. Choices made at the cognitive architecture level can significantly impact FMware performance, either directly or through their influence on the classical software architecture. Below are some critical considerations:
@@
Picking more powerful FMs within a simple cognitive architecture versus simpler FMs within a more complex cognitive architecture: FMware designers face a unique performance dilemma – they must choose between a simple cognitive architecture with fewer, larger, and more capable FMs (incurring high inference costs per request) versus a more complex cognitive architecture that combines multiple FMs (lowering inference costs per request, but involving many more inferences). The composition of multiple FMs introduces latency and performance challenges that go beyond those encountered in traditional software.
@@
The inference costs of FMs vary significantly, with some FM inferences being 10 times more expensive than others [33]. Additionally, the cost of each token generated from a single prompt is not constant. The first token incurs a much higher cost than subsequent tokens due to the need for a KV cache fill, while following tokens reuse this cache to respond faster [17, 34]. These cost dynamics are further complicated by the introduction of the new OpenAI o1 FM, which requires more reasoning time before responding, dramatically increasing the first token’s cost  [35].
@@
Cognitive architecture choices range from leveraging a single FM for basic interactions to complex architectures proposed in multi-agent systems  [36, 37]. Studies and our experiences indicate that smaller FMs within a more complex cognitive architecture can achieve similar, if not better, improvements in FMware quality  [5, 2]. However, increasing cognitive architecture complexity may result in higher latency for end-users (e.g., agents powered by weaker FMs debating each other versus a single prompt to a larger FM  [38]).
@@
Chen et al. [39] demonstrated a balanced approach to FM algorithm design, considering both error reduction and cost minimization metrics. They tuned the parallel decomposition granularity as a hyperparameter, systematically balancing competing error and performance objectives.
@@
Pipelining the execution of cognitive code as it is being generated versus waiting for the full generation and verification of such code: FMware often generates a significant portion of their source code on the fly, either by prompting an FM or through interactions with one or more AI agents. For instance, an FM might be queried to define the necessary steps (i.e., create a plan), which are then executed using FM-powered components or traditional software components. Developers can either wait for the entire set of auto-generated instructions to be completed and verified before executing them  [40, 41], or start pipelining the execution, risking the need to undo steps if the overall plan is later found to be inappropriate  [42].
@@
The addition of semantic caching throughout the cognitive architecture: Semantic caching minimizes FM or AI component inference calls by identifying similar requests or those likely to generate previously produced content. These caches are vital in optimizing the performance of FMware by reducing redundant processing and lowering latency. However, designing caching mechanisms for FMware components remains ad-hoc, lacking best practices or techniques to help architects assess the ROI of adding such caches.
@@
III-B Challenge 2: The complexity of creating token-efficient communication language between the AI components of FMware
@@
Deciding the communication language: Different natural languages require varying amounts of tokens to express the same information semantically (language efficiency and density). This disparity in word-to-token ratios across languages can significantly impact meeting performance requirements [43]. For instance, Hindi requires eight times as many tokens as English to convey the same information [44]. This discrepancy results in longer processing times and varying performance based on the communication language used across the AI components of an FMware. API-based hosted models suffer from increased costs and longer response times with more tokens, while self-hosted models allow for language-specific fine-tuning to mitigate performance impacts.
@@
Prior studies have sought to address the impacts of the varying word-to-token ratios. Nag et al. [45] found that low-resource languages (LRLs) cost more than high-resource languages (HRLs) due to producing more tokens for the same content. They proposed using translation to reduce the token count processed by LRLs. However, adding translation as an intermediate step introduces drawbacks, such as increased processing time, which can affect FMware’s ability to meet SLA requirements. In a prior multilingual FMware project [2], we translated requests to English, used English for internal cognitive communication, then translated responses back. This approach improved performance despite the additional translation costs and aided developers who were not fluent in all supported languages in debugging the FMware.
@@
Further research is needed to design multilingual applications that maintain consistent end-to-end SLAs despite token count disparities. Possible approaches include assigning powerful GPUs for LRLs to speed up processing, adopting Nag et al.’s [45] translation step, and exploring prompt-compression techniques [46] to reduce token counts while retaining essential information. Fine-tuning FMs can also help them better understand and process the unique characteristics of specific domains, mitigating performance impacts due to token disparities.
@@
Defining the communication format: Once the communication language is decided, defining the communication format becomes crucial. JSON is a popular format, fine-tuned by many FMs for its structured, readable, and easily parsable nature [47]. However, JSON often uses more tokens than necessary to convey simple information. Alternatively, a more compact format like YAML, which is less verbose, may use fewer tokens for the same message. Using a less verbose format can make the process more efficient by reducing the time needed to process prompts and generate responses. However, format selection requires careful consideration due to FM biases towards specific output formats. Long et al. [48] found that most FMs generate correctly formatted JSON responses more reliably than YAML, likely because JSON is more prevalent in model training data. LinkedIn’s shift from JSON to YAML for optimizing communication format also highlights these considerations [49].
@@
Similar to human languages, grammar complexity affects performance. For instance, using verbose grammar with complex wording can negatively impact performance. Tam et al. [50] found YAML to be a more cost-effective format for models like GPT-3.5-Turbo compared to JSON, with both text and YAML formats showing lower token generation costs than JSON.
@@
Existing research attempts to leverage less verbose formats for higher performance. Bottaro and Ramgopal [49] noted that after switching to YAML for its brevity, FMs produced invalid output formats 10% of the time. Some studies explored the implications of enforcing constraints on output structure. For example, Kellner et al. [51] observed performance degradation with structural constraints but proposed speculative decoding to minimize overhead and speed up generation.
@@
Chen et al. [52] found that using structured formats like JSON objects, tables, and markdown enhances clarity, accuracy, and reasoning efficiency in FMs, simplifying cognitive architecture and reducing error-handling needs. They proposed AutoForm, an automatic method to select and use the most suitable communication format for a task. Kurt [53] suggested using finite-state machines and regular expressions to enforce structural constraints, improving structured output generation. Nonetheless, further work is needed to reduce invalid outputs across different schemas.
@@
To mitigate these costs, some solutions integrate classical robust-parsing techniques on the communication channels. Instead of spending too many tokens to ensure the quality of the communication protocol, these techniques can help parse FM responses more efficiently. A practical example of this is documented by Bottaro and Ramgopal [49], where they used a classic, CPU-powered robust YAML parser to detect errors in communication. This method helps maintain a low error rate (0.01%), while also saving GPU jobs for more intensive tasks. By offloading the parsing to a CPU, they reduce the need for additional tokens in the prompt, leading to more efficient processing. Another example from Strong [54] proposes a multi step pipeline approach to mitigate the correctness of the output structure where the output structuring step is separated out from actual model reasoning step to produce the correct structured output finally. But the proposed approach uses two inference calls which would increase both cost as well as latency.
Offloading the output parsing and structure formatting to less costly CPU based solutions is a first step towards addressing this challenge. For instance, the output structuring step from Strong’s work [54] can be offloaded to a CPU before sending the result downstream. On the other hand, innovative decoding approaches (such as the one proposed by Beurer-Kellne et al. [51]) which minimizes the performance overhead introduced with output structured generation, is another direction.
@@
A practical implementation of this concept is seen in the work of dottxt team [53]. They proposed the Coalescence framework to speed up the inference by five times with their structured generation that skips unnecessarily calls to FMs leveraging the known structure of the responses, only generating the variable parts that change. This work proposes an efficient guided text generation technique using finite-state machines and regular expressions to enforce structural constraints, significantly reducing computational costs and enhancing output quality while being model-agnostic.
III-C Challenge 3: The complexity of performance tuning and optimization of FMware
Performance tuning and optimization in FMware requires a deep understanding of performance bottlenecks. The core of FMware is the inference of FMs. While many techniques focus on optimizing models [21], efficiently serving FMs is only the beginning. FMware involves interactions among multiple FMs and software components within a cognitive architecture, similar to classical software architecture, where each component has distinct resource demands. This leads to numerous configuration knobs, further complicated by heterogeneous hardware. Additionally, FMware might evolve continuously by itself as its agents perform self-exploration, compared to regular software which is static. Optimizing live FMware is akin to hitting a moving target. Hence, we categorize the challenges into three dimensions as described below:
Complex model-level optimization: Techniques in FMware focus on enhancing hardware utilization, reducing latency, and maximizing throughput during the inference process. Existing FMs mostly rely on decoder-only transformer-based architectures. The inference process for these models has been described in detail in Section II. Many optimization techniques have been proposed, including model architecture redesigns (e.g., multi-query attention) and model compression strategies (e.g., knowledge distillation, quantization) [22]. For a more in-depth understanding, the reader can refer to existing surveys [22, 21]. In this section, we focus on the techniques that directly impact developers of FMware, where they interact with models through prompting.
@@ III-C Challenge 3: The complexity of performance tuning and optimization of FMware
Currently, prompt tuning relies heavily on manual and empirical methods. Developers frequently engage in trial-and-error approaches to refine prompts and find the optimal parameters. For instance, Chen et al. [39] reasoned about the pros and cons of task decomposition for LLM-based applications, where each task formats a prompt based on its input and feeds it into an LLM. They studied parallel decomposition to guide developers in achieving the expected accuracy or efficiency. To boost model inference performance, Kurt [53] leveraged finite state machines and regular expressions to represent deterministic structures in the output, allowing the model to skip over predictable parts of the structure, thus substantially reducing generation latency. Streaming techniques have also been proposed to enhance performance by overlapping the output generation and input for the next model. For example, Bottaro and Ramgopal [49] proposed streaming the application pipeline so that downstream calls can be invoked as soon as they are ready, without waiting for the complete response. Additionally, Santhanam et al. [55] introduced ALTO, an FM serving system for streaming AI pipelines, demonstrating improved throughput and tail latency by streaming intermediate outputs to downstream tasks.
@@
Excessive amount of performance configuration knobs: When dealing with FMware such as those illustrated by OPEA [56] in Figure 1, the configuration landscape becomes significantly complex. As shown, a performance engineer must consider multiple optimization opportunities, including different cognitive architectures, prompt designs, base model selections, model quantization decisions, fine-tuning processes, and communication protocol adjustments. These various aspects, from data ingestion to LLM inference and retrieval, illustrate the intricate array of configuration knobs required for optimizing the application-level performance. We describe three most prominent aspects as follows:
Figure 1: A flowchart of the RAG-based LLM pipeline from OPEA [56], with added annotations on the right panel to highlight key tuning parameters and decision points, illustrating the complexities involved in FMware development
@@
Many solutions have been proposed to address these challenges, each tackling a specific aspect of FMware optimization. Maurya et al. [57] proposed SelectLLM, a framework that analyzes user prompts and selects the most appropriate models at runtime. This approach enables developers to maintain response quality while reducing computational costs, thereby improving the efficiency of model selection. Similarly, Shekhar et al. [58] introduced QC-Opt, a Quality-aware Cost Optimized LLM routing engine and framework. QC-Opt optimizes both the choice of LLM and input token count at runtime to minimize costs while maintaining output quality. This helps developers navigate trade-offs between quality and cost, providing flexibility in selecting models that best fit their requirements. Gong et al. [59] developed a benchmarking toolkit to evaluate various quantization strategies and parameter configurations, providing insights that can guide developers in making decisions on pruning and optimizing models across different deployment scenarios.
To find suitable acceleration inference engines, while some articles provide high-level discussions and benchmarks for different engines  [60], there is still a lack of clear guidelines and standards to make informed selections. Further research and benchmarking are needed for specific scenarios. For instance, Xiao et al. [61] investigated the pros and cons of MLC-LLM and Llama.cpp in mobile environments, using mobile-sensitive metrics such as battery power consumption, latency, and memory bottleneck. These insights are essential for tailoring LLM deployments to mobile devices where resource constraints are more stringent.
To track the complexities of FMware tuning in a holistic perspective, Sun et al. [62] provided a multi-objective benchmarking toolkit, CEBench, that focuses on balancing expenditure and effectiveness for LLM deployments. By allowing easy modifications through configuration files, CEBench supports holistic decision-making across the entire software stack, enabling developers to optimize resource allocation, cost, and performance in an integrated manner. Papaioannou et al. [63] proposed a holistic approach to tuning LLM applications by addressing the complexities introduced by diverse workloads and real-world conditions. They noticed that most LLM applications, which often rely on synthetic datasets, may not account for the variability in input sizes and task demands found in practical applications. Their analysis, which includes different workload types and memory configurations, helps identify key performance bottlenecks and optimization opportunities. By providing a framework that considers a wide range of use cases, their research guides developers in making more informed decisions to enhance FMware efficiency across various scenarios.
@@
Evolving and moving target: Unlike traditional software, FMware is live software that keeps evolving. Each round of execution of an agent might lead to an adjustment of the whole system. To make things worse, agents can self-evolve, making benchmarking much harder than traditional software. As a result, performance issues might be difficult to reproduce due to: (a) the probabilistic nature of the model inference process with token sampling, and (b) the evolving nature of FMware through Data Flywheel [5] or self-exploration.
@@
To address the challenges of reproducibility, a common trick, also suggested by OpenAI [64], is using settings like a low temperature parameter value to ensure more consistent inference outputs. Additionally, setting seeds beforehand, as suggested by PyTorch’s reproducibility guidelines [65], can further help achieve consistent behaviour across repeated inferences. These methods help standardize model behaviour, simplifying the identification of performance bottlenecks in FMware. However, such reproducibility measures restrict the model’s ability to autonomously explore and optimize.
@@
The optimization process of FM and FM-powered agents should evolve from manually-tuning to a Data Flywheel-driven continuous self improving system. Firstly, we must continuously monitor and test if the accuracy of FMware drops through online feedback. Machmouchi and Gupta [66] proposed a comprehensive framework for evaluating LLMs, emphasizing the need for continuous testing with real-time user feedback. They highlighted the importance of segmenting user data to better capture the output quality of FMware. Such a framework would help us understand whether FMs need to be evolved.
@@
The evolution of FMware requires developers to make decisions about how frequently to update the models. One possible solution is fine-tuning the models based on newly generated datasets. Alternatively, developers can update prompting and post-processing techniques to enhance output quality and save costs by not retraining the model. Developers must weigh the cost of fine-tuning as a significant investment against the use of efficient prompting and post-processing techniques applied at the individual request level. Xia et al. [67] proposed a profiling tool to help developers estimate the cost of LLM fine-tuning on GPU clusters, which aids in planning the frequency of fine-tuning based on cost considerations. The decision between these strategies often hinges on factors like user request volume and the desired performance level. In low-volume scenarios, investing in advanced prompting techniques can be more cost-effective because each request can afford additional computational time. Conversely, in high request volume environments, model fine-tuning becomes more beneficial as it allows simpler prompts, thereby reducing the computational overhead for each request.

### v2

@@ 3.2. Challenge 1: The complexity of creating high-performance cognitive architectures
The first step in developing FMware is to create an appropriate cognitive architecture. A cognitive architecture defines how different AI components interact and reason together to achieve desired outcomes. Cognitive architectures for FMware range from simple single-FM pipelines to multi-FM orchestration such as RAG pipelines, tool-augmented chains, and workflow- or DAG-based prompt graphs. More advanced forms include autonomous agent architectures, where FM modules plan, invoke tools, and collaborate dynamically to execute multi-step reasoning and actions. The FMware cognitive architecture complements the classical software architecture, detailing how AI reasoning and results are delivered through traditional software components like regular and/or vector databases. Choices made at the cognitive architecture level can significantly impact FMware performance, either directly or through their influence on the classical software architecture. Below are some critical considerations:
@@
Picking more powerful FMs within a simple cognitive architecture versus simpler FMs within a more complex cognitive architecture: FMware designers face a unique performance dilemma – they must choose between a simple cognitive architecture with fewer, larger, and more capable FMs (incurring high inference costs per request) versus a more complex cognitive architecture that combines multiple FMs (lowering inference costs per request, but involving many more inferences). We highlight that the choice of the cognitive architecture depends on the specific use case; however, generally speaking the composition of multiple FMs introduces latency and performance challenges that go beyond those encountered in traditional software.
@@
The inference costs of FMs vary significantly, with some FM inferences being 10 times more expensive than others (Arc53, 2024). Additionally, the cost of each token generated from a single prompt is not constant. The first token incurs a much higher cost than subsequent tokens due to the need for a KV cache fill, while following tokens reuse this cache to respond faster (Agrawal et al., 2023; Yuan et al., 2024b). These cost dynamics are further complicated by the introduction of the new OpenAI o1 FM, which requires more reasoning time before responding, dramatically increasing the first token’s cost  (OpenAI, 2024a).
@@
Cognitive architecture choices range from leveraging a single FM for basic interactions to complex architectures proposed in multi-agent systems  (Sumers et al., arXiv preprint arXiv:2410.03613, 2024; Wu et al., arXiv preprint arXiv:2308.08155, 2023). Studies and our experiences indicate that smaller FMs within a more complex cognitive architecture can achieve similar, if not better, improvements in FMware quality  (Yan et al., 2024a; Hassan et al., 2024). However, increasing cognitive architecture complexity may result in higher latency for end-users (e.g., agents powered by weaker FMs debating each other versus a single prompt to a larger FM  (Chan et al., arXiv preprint arXiv:2308.07201, 2023)).
@@
Chen et al. (Chen et al., 2024b) demonstrated a balanced approach to FM algorithm design, considering both error reduction and cost minimization metrics. They tuned the parallel decomposition granularity as a hyperparameter, systematically balancing competing error and performance objectives.
@@
Pipelining the execution of cognitive code as it is being generated versus waiting for the full generation and verification of such code: FMware often generates a significant portion of their source code on the fly, either by prompting an FM or through interactions with one or more AI agents. For instance, an FM might be queried to define the necessary steps (i.e., create a plan), which are then executed using FM-powered components or traditional software components. Developers can either wait for the entire set of auto-generated instructions to be completed and verified before executing them  (Zhou et al., 2024b; Wang et al., 2023), or start pipelining the execution, risking the need to undo steps if the overall plan is later found to be inappropriate  (Chen and Li, 2024).
@@
The addition of semantic caching throughout the cognitive architecture: A semantic cache provides a mechanism to use prior user prompts and FM completions to address similar user prompts by using vector similarity search (Corporation, 2025). Semantic caching minimizes FM or AI component inference calls by identifying similar requests or those likely to generate previously produced content. These caches are vital in optimizing the performance of FMware by reducing redundant processing and lowering latency. For example, Regmi et al. (Regmi and Pun, 2024) demonstrate that caching query embeddings in a GPT Semantic Cache can reduce LLM API calls by up to 68.8% and significantly lower inference latency. Similarly, Gill et al. (Gill et al., 2024) shows how retrieving past responses for semantically similar queries avoids repeated model invocation and reduces the latency and costs. However, designing caching mechanisms for FMware components remains ad-hoc, lacking best practices or techniques to help architects assess the ROI of adding such caches.
@@
3.3. Challenge 2: The complexity of creating token-efficient communication language between the AI components of FMware
@@
Deciding the communication language: Different natural languages require varying amounts of tokens to express the same information semantically (language efficiency and density). This disparity in word-to-token ratios across languages can significantly impact meeting performance requirements (Petrov et al., 2024). For instance, Hindi requires eight times as many tokens as English to convey the same information (Reddit, 2024). This discrepancy results in longer processing times and varying performance based on the communication language used across the AI components of an FMware. API-based hosted models suffer from increased costs and longer response times with more tokens, while self-hosted models allow for language-specific fine-tuning to mitigate performance impacts.
@@
Prior studies have sought to address the impacts of the varying word-to-token ratios. Nag et al. (Nag et al., 2024) found that low-resource languages (LRLs) cost more than high-resource languages (HRLs) due to producing more tokens for the same content. They proposed using translation to reduce the token count processed by LRLs. However, adding translation as an intermediate step introduces drawbacks, such as increased processing time, which can affect FMware’s ability to meet SLA requirements. In a prior multilingual FMware project (Hassan et al., 2024), we translated requests to English, used English for internal cognitive communication, then translated responses back. This approach improved performance despite the additional translation costs and aided developers who were not fluent in all supported languages in debugging the FMware.
@@
Further research is needed to design multilingual applications that maintain consistent end-to-end SLAs despite token count disparities. Possible approaches include assigning powerful GPUs for LRLs to speed up processing, adopting Nag et al.’s (Nag et al., 2024) translation step, and exploring prompt-compression techniques (Jiang et al., 2023) to reduce token counts while retaining essential information. Fine-tuning FMs can also help them better understand and process the unique characteristics of specific domains, mitigating performance impacts due to token disparities.
@@
Defining the communication format: Once the communication language is decided, defining the communication format becomes crucial. JSON is a popular format, fine-tuned by many FMs for its structured, readable, and easily parsable nature (Liu et al., 2024a). However, JSON often uses more tokens than necessary to convey simple information. Alternatively, a more compact format like YAML, which is less verbose, may use fewer tokens for the same message. Using a less verbose format can make the process more efficient by reducing the time needed to process prompts and generate responses. However, format selection requires careful consideration due to FM biases towards specific output formats. Long et al. (Long et al., 2024) found that most FMs generate correctly formatted JSON responses more reliably than YAML, likely because JSON is more prevalent in model training data. LinkedIn’s shift from JSON to YAML for optimizing communication format also highlights these considerations (Bottaro and Ramgopal, 2024).
@@
Similar to human languages, grammar complexity affects performance. For instance, using verbose grammar with complex wording can negatively impact performance. Tam et al. (Tam et al., 2024) found YAML to be a more cost-effective format for models like GPT-3.5-Turbo compared to JSON, with both text and YAML formats showing lower token generation costs than JSON.
@@
Existing research attempts to leverage less verbose formats for higher performance. Bottaro and Ramgopal (Bottaro and Ramgopal, 2024) noted that after switching to YAML for its brevity, FMs produced invalid output formats 10% of the time. Some studies explored the implications of enforcing constraints on output structure. For example, Kellner et al. (Beurer-Kellner et al., 2024) observed performance degradation with structural constraints but proposed speculative decoding to minimize overhead and speed up generation.
@@
Chen et al. (Chen et al., 2024a) found that using structured formats like JSON objects, tables, and markdown enhances clarity, accuracy, and reasoning efficiency in FMs, simplifying cognitive architecture and reducing error-handling needs. They proposed AutoForm, an automatic method to select and use the most suitable communication format for a task. Kurt (Kurt, 2024) suggested using finite-state machines and regular expressions to enforce structural constraints, improving structured output generation. Nonetheless, further work is needed to reduce invalid outputs across different schemas.
@@
To mitigate these costs, some solutions integrate classical robust-parsing techniques on the communication channels. Instead of spending too many tokens to ensure the quality of the communication protocol, these techniques can help parse FM responses more efficiently. A practical example of this is documented by Bottaro and Ramgopal (Bottaro and Ramgopal, 2024), where they used a classic, CPU-powered robust YAML parser to detect errors in communication. This method helps maintain a low error rate (0.01%), while also saving GPU jobs for more intensive tasks. By offloading the parsing to a CPU, they reduce the need for additional tokens in the prompt, leading to more efficient processing. Another example from Strong (Strong, 2024) proposes a multi step pipeline approach to mitigate the correctness of the output structure where the output structuring step is separated out from actual model reasoning step to produce the correct structured output finally. But the proposed approach uses two inference calls which would increase both cost as well as latency.
Offloading the output parsing and structure formatting to less costly CPU based solutions is a first step towards addressing this challenge. For instance, the output structuring step from Strong’s work (Strong, 2024) can be offloaded to a CPU before sending the result downstream. On the other hand, innovative decoding approaches (such as the one proposed by Beurer-Kellne et al. (Beurer-Kellner et al., 2024)) which minimizes the performance overhead introduced with output structured generation, is another direction.
@@
A practical implementation of this concept is seen in the work of dottxt team (Kurt, 2024). They proposed the Coalescence framework to speed up the inference by five times with their structured generation that skips unnecessarily calls to FMs leveraging the known structure of the responses, only generating the variable parts that change. This work proposes an efficient guided text generation technique using finite-state machines and regular expressions to enforce structural constraints, significantly reducing computational costs and enhancing output quality while being model-agnostic.
3.4. Challenge 3: The complexity of performance tuning and optimization of FMware
Performance tuning and optimization in FMware requires a deep understanding of performance bottlenecks. The core of FMware is the inference of FMs. While many techniques focus on optimizing models (Zhou et al., 2024c), efficiently serving FMs is only the beginning. FMware involves interactions among multiple FMs and software components within a cognitive architecture, similar to classical software architecture, where each component has distinct resource demands. This leads to numerous configuration knobs, further complicated by heterogeneous hardware. Additionally, FMware might evolve continuously by itself as its agents perform self-exploration, compared to regular software which is static. Optimizing live FMware is akin to hitting a moving target. Hence, we categorize the challenges into three dimensions as described below:
Complex model-level optimization: Techniques in FMware focus on enhancing hardware utilization, reducing latency, and maximizing throughput during the inference process. Existing FMs mostly rely on decoder-only transformer-based architectures. The inference process for these models has been described in detail in Section 2. Many optimization techniques have been proposed, including model architecture redesigns (e.g., multi-query attention) and model compression strategies (e.g., knowledge distillation, quantization) (Wang et al., 2024c). For a more in-depth understanding, the reader can refer to existing surveys (Wang et al., 2024c; Zhou et al., 2024c). In this section, we focus on the techniques that directly impact developers of FMware, where they interact with models through prompting.
@@ 3.4. Challenge 3: The complexity of performance tuning and optimization of FMware
Currently, prompt tuning relies heavily on manual and empirical methods. Developers frequently engage in trial-and-error approaches to refine prompts and find the optimal parameters. For instance, Chen et al. (Chen et al., 2024b) reasoned about the pros and cons of task decomposition for LLM-based applications, where each task formats a prompt based on its input and feeds it into an LLM. They studied parallel decomposition to guide developers in achieving the expected accuracy or efficiency. To boost model inference performance, Kurt (Kurt, 2024) leveraged finite state machines and regular expressions to represent deterministic structures in the output, allowing the model to skip over predictable parts of the structure, thus substantially reducing generation latency. Streaming techniques have also been proposed to enhance performance by overlapping the output generation and input for the next model. For example, Bottaro and Ramgopal (Bottaro and Ramgopal, 2024) proposed streaming the application pipeline so that downstream calls can be invoked as soon as they are ready, without waiting for the complete response. Additionally, Santhanam et al. (Santhanam et al., 2024) introduced ALTO, an FM serving system for streaming AI pipelines, demonstrating improved throughput and tail latency by streaming intermediate outputs to downstream tasks.
@@
Excessive amount of performance configuration knobs: When dealing with FMware such as those illustrated by OPEA (Open Exploration of AI (OPEA), 2024) in Figure 2, the configuration landscape becomes significantly complex. As shown, a performance engineer must consider multiple optimization opportunities, including different cognitive architectures, prompt designs, base model selections, model quantization decisions, fine-tuning processes, and communication protocol adjustments. These various aspects, from data ingestion to LLM inference and retrieval, illustrate the intricate array of configuration knobs required for optimizing the application-level performance. We describe three most prominent aspects as follows:
Figure 2. A flowchart of the RAG-based LLM pipeline from OPEA (Open Exploration of AI (OPEA), 2024), with added annotations on the right panel to highlight key tuning parameters and decision points, illustrating the complexities involved in FMware development
@@
Many solutions have been proposed to address these challenges, each tackling a specific aspect of FMware optimization. Maurya et al. (Maurya et al., 2024) proposed SelectLLM, a framework that analyzes user prompts and selects the most appropriate models at runtime. This approach enables developers to maintain response quality while reducing computational costs, thereby improving the efficiency of model selection. Similarly, Shekhar et al. (Shekhar et al., 2024) introduced QC-Opt, a Quality-aware Cost Optimized LLM routing engine and framework. QC-Opt optimizes both the choice of LLM and input token count at runtime to minimize costs while maintaining output quality. This helps developers navigate trade-offs between quality and cost, providing flexibility in selecting models that best fit their requirements. Gong et al. (Gong et al., 2024) developed a benchmarking toolkit to evaluate various quantization strategies and parameter configurations, providing insights that can guide developers in making decisions on pruning and optimizing models across different deployment scenarios.
To find suitable acceleration inference engines, while some articles provide high-level discussions and benchmarks for different engines  (Zhou et al., 2024a), there is still a lack of clear guidelines and standards to make informed selections. Further research and benchmarking are needed for specific scenarios. For instance, Xiao et al. (Xiao et al., 2024) investigated the pros and cons of MLC-LLM and Llama.cpp in mobile environments, using mobile-sensitive metrics such as battery power consumption, latency, and memory bottleneck. These insights are essential for tailoring LLM deployments to mobile devices where resource constraints are more stringent.
To track the complexities of FMware tuning in a holistic perspective, Sun et al. (Sun et al., 2024b) provided a multi-objective benchmarking toolkit, CEBench, that focuses on balancing expenditure and effectiveness for LLM deployments. By allowing easy modifications through configuration files, CEBench supports holistic decision-making across the entire software stack, enabling developers to optimize resource allocation, cost, and performance in an integrated manner. Papaioannou et al. (Papaioannou and Doudali, 2024) proposed a holistic approach to tuning LLM applications by addressing the complexities introduced by diverse workloads and real-world conditions. They noticed that most LLM applications, which often rely on synthetic datasets, may not account for the variability in input sizes and task demands found in practical applications. Their analysis, which includes different workload types and memory configurations, helps identify key performance bottlenecks and optimization opportunities. By providing a framework that considers a wide range of use cases, their research guides developers in making more informed decisions to enhance FMware efficiency across various scenarios.
@@
Evolving and moving target: Unlike traditional software, FMware is live software that keeps evolving. Each round of execution of an agent might lead to an adjustment of the whole system. To make things worse, agents can self-evolve, making benchmarking much harder than traditional software. As a result, performance issues might be difficult to reproduce due to: (a) the probabilistic nature of the model inference process with token sampling, and (b) the evolving nature of FMware through Data Flywheel (Yan et al., 2024a) or self-exploration.
@@
To address the challenges of reproducibility, a common trick, also suggested by OpenAI (OpenAI, 2024b), is using settings like a low temperature parameter value to ensure more consistent inference outputs. Additionally, setting seeds beforehand, as suggested by PyTorch’s reproducibility guidelines (PyTorch, 2024), can further help achieve consistent behaviour across repeated inferences. These methods help standardize model behaviour, simplifying the identification of performance bottlenecks in FMware. However, such reproducibility measures restrict the model’s ability to autonomously explore and optimize.
@@
The optimization process of FM and FM-powered agents should evolve from manually-tuning to a Data Flywheel-driven continuous self improving system. Firstly, we must continuously monitor and test if the accuracy of FMware drops through online feedback. Machmouchi and Gupta (Machmouchi and Gupta, 2023) proposed a comprehensive framework for evaluating LLMs, emphasizing the need for continuous testing with real-time user feedback. They highlighted the importance of segmenting user data to better capture the output quality of FMware. Such a framework would help us understand whether FMs need to be evolved.
@@
The evolution of FMware requires developers to make decisions about how frequently to update the models. One possible solution is fine-tuning the models based on newly generated datasets. Alternatively, developers can update prompting and post-processing techniques to enhance output quality and save costs by not retraining the model. Developers must weigh the cost of fine-tuning as a significant investment against the use of efficient prompting and post-processing techniques applied at the individual request level. Xia et al. (Xia et al., 2024) proposed a profiling tool to help developers estimate the cost of LLM fine-tuning on GPU clusters, which aids in planning the frequency of fine-tuning based on cost considerations. The decision between these strategies often hinges on factors like user request volume and the desired performance level. In low-volume scenarios, investing in advanced prompting techniques can be more cost-effective because each request can afford additional computational time. Conversely, in high request volume environments, model fine-tuning becomes more beneficial as it allows simpler prompts, thereby reducing the computational overhead for each request.